jigglypuff314 (jigglypuff314):

If increasing 15 by n% is equivalent to decreasing 75 by n%, then what is 6% of n?

OpenStudy (raden):

first, find n. 15 + n%(15) = 75 - n%(75) solve for n = ?

jigglypuff314 (jigglypuff314):

would it be n% = 6/9 ?

OpenStudy (raden):

yup, that's correct

jigglypuff314 (jigglypuff314):

okay, I get it now :) thank you!

OpenStudy (raden):

really, that's cool. i got 4 as the answer, same with you ?

jigglypuff314 (jigglypuff314):

yep :D

OpenStudy (raden):

ah, okay... :D